# Influence of adversarial training on super-resolution turbulence reconstruction

### v4

@@ Abstract
Supervised super-resolution deep convolutional neural networks (CNNs) have gained significant attention for their potential in reconstructing velocity and scalar fields in turbulent flows. Despite their popularity, CNNs currently lack the ability to accurately produce high-frequency and small-scale features, and tests of their generalizability to out-of-sample flows are not widespread. Generative adversarial networks (GANs), which consist of two distinct neural networks (NNs), a generator and discriminator, are a promising alternative, allowing for both semi-supervised and unsupervised training. The difference in the flow fields produced by these two NN architectures has not been thoroughly investigated, and a comprehensive understanding of the discriminator’s role has yet to be developed. This study assesses the effectiveness of the unsupervised adversarial training in GANs for turbulence reconstruction in forced homogeneous isotropic turbulence. GAN-based architectures are found to outperform supervised CNNs for turbulent flow reconstruction for in-sample cases. The reconstruction accuracy of both architectures diminishes for out-of-sample cases, though the GAN’s discriminator network significantly improves the generator’s out-of-sample robustness using either an additional unsupervised training step with large eddy simulation input fields and a dynamic selection of the most suitable upsampling factor. These enhance the generator’s ability to reconstruct small-scale gradients, turbulence intermittency, and velocity-gradient probability density functions. Conversely, the supervised super-resolution CNN network lacks the capability to reconstruct these statistics. The extrapolation capability of the GAN-based model is demonstrated for out-of-sample flows at higher Reynolds numbers. Based on these findings, incorporating discriminator-based training is recommended to enhance the reconstruction capability of super-resolution CNNs.
@@ 1 Introduction
To address the drawbacks of the supervised CNNs training strategy for SR, generative adversarial networks (GANs) have been employed [13] that, unlike CNNs, do not rely solely on user-defined objective functions for training. Instead, GANs’ super-resolving generator networks, based on CNNs, are coupled with discriminator networks that compete in semi-supervised learning. In GAN-based SR, the generator produces higher-resolution versions of low-resolution fields (e.g., filtered DNS or LES), while the discriminator network attempts to differentiate between the generated high-resolution field and the real, “ground truth” high-resolution field (e.g., the DNS itself). During training, the generator learns to produce samples that are indistinguishable from genuine high-resolution data. The discriminator in turn learns to judge the authenticity of the samples. As a result, both networks improve each other during the adversarial training process.
@@
While numerous studies have recognized the superior performance of GANs for super-resolution reconstruction over conventional supervised CNNs [4, 26, 17, 27], the underlying reasons for this advantage have not been thoroughly explored. Possibilities include the flexibility in the training strategy (supervised and partially unsupervised), the use of adversarial (unsupervised) training leveraging unpaired training data, and the exploitation of generative models to model distributions, thereby improving generalization to out-of-sample data. GANs are not without drawbacks and may exhibit training instability and mode collapse, which could hinder convergence of the generator [28]. Given these and the computational cost of training GANs, a critical evaluation of the need for SR-GANs and the influence of the discriminator on generalizability to both in- and out-of-sample flows with respect to different kernels and sizes and different flow configurations, is necessary. Moreover, the performance of GAN-based models to reconstruct spatial features and correlations, including gradients, intermittency, and probability density function tails, has not been thoroughly examined.
This study investigates and quantifies the role of the GAN discriminator in reconstructing turbulent fields while considering the benefits and limitations of the GAN approach. To evaluate the effectiveness of adversarial training, a supervised deep-CNN model (TSResNet; Sec. 3.1) is directly compared to a semi-supervised/unsupervised GAN (PIESRGAN; Sec. 3.2). The same CNN architecture, physics-inspired loss functions, and training data are used for the TSResNet and PIESRGAN models. Hence, the sole difference between the two lies in the training strategy. In particular, the ability to accurately reconstruct small-scale structures of turbulence, consistent with the fundamental theory of turbulence, is investigated. Furthermore, while most studies have focused on in-sample reconstruction within a narrow range of flow conditions, we explore both models’ generalizability to out-of-sample filtered fields for different filter sizes, filter kernels, and Reynolds numbers. Our aim is to develop an understanding of the consistency of the reconstructed fields with the original flow while avoiding the generation of spurious and/or unphysical fields and preserving the fundamental turbulence characteristics. Particular emphasis is placed on the improvements of these metrics brought about by adversarial training.
@@ 2 Datasets and preprocessing
After the lowest $Re$ simulation reached a statistically stationary state, 160 snapshots of the 3D velocity vector ($u_{i}$ = $(\textit{U},\textit{V},\textit{W})^{T}$) were extracted every five eddy-turnover times for training. To obtain low-resolution data, all available snapshots of the instantaneous velocity were filtered using explicit filter kernels of width $\Delta=4\,dx$, corresponding to $8\,\eta$. Our analysis is limited by memory requirements associated with the upsampling factor. As demonstrated later, the filter size considered places the cut-off wavenumber into the inertial range. As the primary goal is to demonstrate the effectiveness of adversarial training for SR reconstructions by comparing different training strategies, the advantage of adversarial contribution is apparent even at these upsampling factors. The training, validation, and testing datasets were composed of high-resolution (DNS) data, and the corresponding low-resolution (filtered-DNS) data. The spatial box, Gaussian, and spectrally sharp filter kernels are defined as
@@
Loading the entire training dataset is impractical given memory limitations. To alleviate this, we extract sub-domains randomly from each snapshot of the full domain. The first 120 snapshots were used for training, while the last 40 frames, equally divided, were used for validation and testing (training/validation ratio = $0.75$/$0.25$). The ongoing validation was then performed on 20 in-sample fields not used for training. The last 20 snapshots were employed to obtain averaged results. The time separation between training/validation samples and testing samples spans more than 10 large-eddy turnover times, thus the two sets of samples should be statistically uncorrelated. Both networks, TSResNet and PIESRGAN (introduced in Sec. 3), were trained using sub-boxes of the original computational domain with a size of $128\,\eta\times 128\,\eta\times 128\,\eta$, corresponding to sub-boxes of $64\times 64\times 64$ and $16\times 16\times 16$ for high- and low-resolution fields, respectively. The size of these sub-boxes was found to be a good compromise between memory requirements and the inherent length scale of the turbulent flow. It is important to emphasize that the dimensions of the sub-boxes were constrained by both $\eta$ and the integral length scale $l_{t}$. Specifically, it is ensured that at least one $l_{t}$ is contained in these sub-boxes. Using those settings, 6400 sub-boxes were employed for the training. During the evaluation of testing samples, the entire test sample is reconstructed continuously, without employing cropped reconstruction boxes as is done for training. The velocity components of high-resolution data and low-resolution data used for training and testing were normalized by the global maximum and minimum of the DNS (i.e., $(U_{\textsc{f-dns}},V_{\textsc{f-dns}},W_{\textsc{f-dns}})\in[0,1]$ and $(U_{\textsc{dns}},V_{\textsc{dns}},W_{\textsc{dns}})\in[0,1]$) to improve the network’s performance [31].
@@
The large dataset and deep convolutional frameworks (implemented in TensorFlow [32]), required training parallelization; this was done using the Horovod library [33]. Calculations were performed on the Jülich DEEP-EST cluster (DEEP-DAM partition) using four nodes, each with one NVIDIA V100 32GB GPU. This gave an overall speed-up close to a factor of four versus single-GPU training. A mini-batch of 8 sub-boxes per GPU and the ADAM optimizer, with an initial learning rate of $10^{-4}$, were chosen due to memory constraints and previous evidence of the success of this combination [9, 17, 34]. “Mixed-precision” training is common in machine learning for its low computational cost [35] but can be detrimental to predictive accuracy in scientific computing. We use single-precision arithmetic for both weight definition and loss computations, following the results of Hrycej et al. [36]. Details are provided in appendix A. Moreover, during the evaluation of the testing samples, CPUs with access to a significantly larger memory pool are employed to overcome memory limitations.
@@
For unsupervised training, LES calculations of forced HIT are conducted using the same DNS solver configuration albeit for coarser grid resolutions (equivalent to the discrete downsampling factor applied to the DNS mesh). The unresolved scales are closed using the dynamic Smagorinsky model [37, 38] (Eq. 11), or their influence is not explicitly modeled (“implicitly modeled” LES). To extrapolate a given model to higher $Re$ (utilizing the Re130 and Re350 datasets for out-of-sample analyses), the filtered input field is rescaled to match the $\Delta/\eta$ ratio used for training. This rescaling procedure guarantees a certain level of generalizability [21].
@@ 3 Neural network architectures and training strategy
Our networks are based on the architecture of Bode et al. [19, 20, 39, 40]. Differing from previous work, our generator network uses additional up-/downsampling layers and dense layers, depicted in Fig. 1; this is inspired by the original ESRGAN architecture [16] and adapted for small-scale turbulence reconstruction [22]. Along with the definition of the two architectures, the training approaches are described next. In order to enhance the reproducibility of this study and provide additional clarity on technical aspects of the architectures, access to our GIT repository will be provided upon request.
@@ 3.1 TSResNet
The TSResNet (Turbulent Super-Resolution Residual Network) is a supervised CNN-type model. It uses only the generator shown in Fig. 1 without the discriminator. The TSResNet is adapted from the original SRResNet [41], which can capture small-scale features when skip-connections are included. The generator relies on three-dimensional convolutional layers with Leaky Rectified Linear Units (LReLUs) as activation functions, as they have higher computational efficiency compared to standard rectified linear units (ReLUs) [32]. The Residual-in-Residual Dense Blocks (RRDBs) include residual connections, a set of densely-connected-layer blocks (three in the present work), and a residual-in-residual structure [16]. Together, these enable super-resolved data to be generated through a very deep network. This capability is essential to facilitate learning complex transformations and is currently one of the state-of-the-art features of super-resolution networks [42]. Two upsampling layers increase the input spatial resolution by a factor of four in each spatial dimension. Each one doubles the dimensions of the input by nearest-neighbor interpolation (replicating adjacent grid points), followed by a convolution layer to improve the approximation. The upsampling branch (the gray section in Fig. 1) is later modified in Sec. 5.2, where additional upsampling layers are added. The total number of trainable parameters in the TSResNet is approximately $18$ million.
@@ 3.2 PIESRGAN
The training of the GAN architecture is challenging, as the generator and discriminator networks are trained to compete against each other. Finding a convergence point is one of the main challenges of GAN training, and training can suffer oscillations and destabilization in the model’s trainable parameters [13]. To cope with these, the generator is pre-trained in a fully-supervised manner (like the TSResNet) by using the original generator loss function proposed in Eq. 6. Subsequently, the pre-trained generator is used to initialize the GAN training. Recent findings demonstrated that the choice of the initial learning rate for both networks is essential for the local convergence of the GAN training, and recommendations from recent literature have been adopted [28, 22].
@@ 3.3 Training strategy
Table 2 lists the architecture, training strategy, and training/testing fields for the specific trained models used in subsequent sections. In the supervised and semi-supervised approaches, both LR and DNS field pairs are accessible, and the network utilizes this paired dataset in an attempt to learn the mapping function between LR fields and to their corresponding ground-truth counterparts. This methodology is viable only when the dataset contains paired data, i.e., both F-DNS and DNS data. In the unsupervised approach, only LR fields (LES data) are available without the corresponding DNS fields, thus learning relies solely on the feedback provided by the discriminator and the physics-informed loss function. The partially supervised approach is a combination of the two strategies, in which F-DNS/DNS pairs and true LES data as well as the discriminator feedback are employed for training.
The TSResNet models were trained for roughly 200 epochs, after which no additional reconstruction improvement was observed, using the Re90 dataset and Eq. 6 as the loss function. Statistical convergence required 37 wall-time hours using four NVIDIA Tesla V100 32GB GPUs. The trained TSResNet model parameters were used to initialize the GAN generators. No GAN discriminator pre-training was applied, as this step has recently been demonstrated to have minimal influence [22]. The adversarial training required an additional 11 wall-time hours using the same hardware, i.e., roughly $30\,\%$ more than the TSResNet training time, over an additional 40 epochs. To counteract the influence of the additional training time performed during the GAN training, the TSResNet models were trained for an additional 40 epochs. However, this additional training time did not impact the final reconstructed field and can be considered superfluous.
Model Training strategy Input training – testing fields: Filter kernels ($\overline{\Delta}/\Delta_{\textsc{dns}}$) Section
@@
TSResNet (spectral kernel) supervised S ($4$) – S ($4$) §4
TSResNet (box kernel) supervised B ($4$) – B ($4$) §4
PIESRGAN (spectral kernel) semi-supervised S ($4$) – S ($4$) §4
PIESRGAN (box kernel) semi-supervised B ($4$) – B ($4$) §4
@@
PIESRGAN (multiple-kernel-trained) semi-supervised B $\land$ S ($4$) – G ($4$) §5.1
PIESRGAN (partially unsupervised) semi-supervised &\&& unsupervised B $\land$ S $\land$ LES ($4$) – G ($4$) §5.1
TSResNet (fixed upsampling) supervised B ($4$) – B ($8$) §5.2
PIESRGAN (fixed upsampling) semi-supervised B ($4$) – B ($8$) §5.2
PIESRGAN ($x\#$) semi-supervised &\&& dynamic upsampling B ($2$land $4$land $8$) – B (##\##) §5.2
@@
TSResNet supervised (Re90) B ($4$) – B ($4$) §6
PIESRGAN semi-supervised (Re90) B ($4$) – B ($4$) §6
Table 2: List of model configurations, training strategies, and training/testing filter kernels and sizes. The filter kernels “B,” “G,” and “S” abbreviate box, Gaussian, and spectrally-sharp kernels. The $\land$ indicates that the combined set of the input fields is employed in the training, while the ##\## indicates the variable upsampling factor.

### v5

@@ Abstract
Supervised super-resolution deep convolutional neural networks (CNNs) have gained significant attention for their potential in reconstructing velocity and scalar fields in turbulent flows. Despite their popularity, CNNs currently lack the ability to accurately produce high-frequency and small-scale features, and tests of their generalizability to out-of-sample flows are not widespread. Generative adversarial networks (GANs), which consist of two distinct neural networks (NNs), a generator and discriminator, are a promising alternative, allowing for both semisupervised and unsupervised training. The difference in the flow fields produced by these two NN architectures has not been thoroughly investigated, and a comprehensive understanding of the discriminator’s role has yet to be developed. This study assesses the effectiveness of the unsupervised adversarial training in GANs for turbulence reconstruction in forced homogeneous isotropic turbulence. GAN-based architectures are found to outperform supervised CNNs for turbulent flow reconstruction for in-sample cases. The reconstruction accuracy of both architectures diminishes for out-of-sample cases, though the GAN’s discriminator network significantly improves the generator’s out-of-sample robustness using either an additional unsupervised training step with large eddy simulation input fields or a dynamic selection of the most suitable upsampling factor. These enhance the generator’s ability to reconstruct small-scale gradients, turbulence intermittency, and velocity-gradient probability density functions. Conversely, the supervised super-resolution CNN network lacks the capability to reconstruct these statistics. The extrapolation capability of the GAN-based model is demonstrated for out-of-sample flows at higher Reynolds numbers. Based on these findings, incorporating discriminator-based training is recommended to enhance the reconstruction capability of super-resolution CNNs.
@@ 1 Introduction
To address the drawbacks of the supervised CNNs training strategy for SR, generative adversarial networks (GANs) have been employed [13] that, unlike CNNs, do not rely solely on user-defined objective functions for training. Instead, GANs’ super-resolving generator networks, based on CNNs, are coupled with discriminator networks that compete in semisupervised learning. In GAN-based SR, the generator produces higher-resolution versions of low-resolution fields (e.g., filtered DNS or LES), while the discriminator network attempts to differentiate between the generated high-resolution field and the real, “ground truth” high-resolution field (e.g., the DNS itself). During training, the generator learns to produce samples that are indistinguishable from genuine high-resolution data. The discriminator in turn learns to judge the authenticity of the samples. As a result, both networks improve each other during the adversarial training process.
@@
While numerous studies have recognized the superior performance of GANs for super-resolution reconstruction over conventional supervised CNNs [4, 26, 17, 27, 28], the underlying reasons for this advantage have not been thoroughly explored. Possibilities include the flexibility in the training strategy (supervised and partially unsupervised), the use of adversarial (unsupervised) training leveraging unpaired training data, and the exploitation of generative models to model distributions, thereby improving generalization to out-of-sample data. GANs are not without drawbacks and may exhibit training instability and mode collapse, which could hinder convergence of the generator [29]. Given these and the computational cost of training GANs, a critical evaluation of the need for SR-GANs and the influence of the discriminator on generalizability to both in- and out-of-sample flows with respect to different kernels and sizes and different flow configurations, is necessary. Moreover, the performance of GAN-based models to reconstruct spatial features and correlations, including gradients, intermittency, and probability density function tails, has not been thoroughly examined.
This study investigates and quantifies the role of the GAN discriminator in reconstructing turbulent fields while considering the benefits and limitations of the GAN approach. To evaluate the effectiveness of adversarial training, a supervised deep-CNN model (TSResNet; Sec. 3.1) is directly compared to a semisupervised/unsupervised GAN (PIESRGAN; Sec. 3.2). The same CNN architecture, physics-inspired loss functions, and training data are used for the TSResNet and PIESRGAN models. Hence, the sole difference between the two lies in the training strategy. In particular, the ability to accurately reconstruct small-scale structures of turbulence, consistent with the fundamental theory of turbulence, is investigated. Furthermore, while most studies have focused on in-sample reconstruction within a narrow range of flow conditions, we explore both models’ generalizability to out-of-sample filtered fields for different filter sizes, filter kernels, and Reynolds numbers. Our aim is to develop an understanding of the consistency of the reconstructed fields with the original flow while avoiding the generation of spurious and/or unphysical fields and preserving the fundamental turbulence characteristics. Particular emphasis is placed on the improvements of these metrics brought about by adversarial training.
@@ 2 Data sets and preprocessing
After the lowest $Re$ simulation reached a statistically stationary state, 160 snapshots of the 3D velocity vector [$u_{i}$ = $(\textit{U},\textit{V},\textit{W})^{T}$] were extracted every five eddy-turnover times for training. To obtain low-resolution data, all available snapshots of the instantaneous velocity were filtered using explicit filter kernels of width $\Delta=4\,dx$, corresponding to $8\,\eta$. Our analysis is limited by memory requirements associated with the upsampling factor. As demonstrated later, the filter size considered places the cut-off wave number into the inertial range. As the primary goal is to demonstrate the effectiveness of adversarial training for SR reconstructions by comparing different training strategies, the advantage of adversarial contribution is apparent even at these upsampling factors. The training, validation, and testing data sets were composed of high-resolution (DNS) data, and the corresponding low-resolution (filtered-DNS) data. The spatial box, Gaussian, and spectrally sharp filter kernels are defined as
@@
Loading the entire training data set is impractical given memory limitations. To alleviate this, we extract sub-domains randomly from each snapshot of the full domain. The first 120 snapshots were used for training, while the last 40 frames, equally divided, were used for validation and testing (training/validation ratio = $0.75$/$0.25$). The ongoing validation was then performed on 20 in-sample fields not used for training. The last 20 snapshots were employed to obtain averaged results. The time separation between training/validation samples and testing samples spans more than 10 large-eddy turnover times, thus the two sets of samples should be statistically uncorrelated. Both networks, TSResNet and PIESRGAN (introduced in Sec. 3), were trained using subboxes of the original computational domain with a size of $128\,\eta\times 128\,\eta\times 128\,\eta$, corresponding to subboxes of $64\times 64\times 64$ and $16\times 16\times 16$ for high- and low-resolution fields, respectively. The size of these subboxes was found to be a good compromise between memory requirements and the inherent length scale of the turbulent flow. It is important to emphasize that the dimensions of the subboxes were constrained by both $\eta$ and the integral length scale $l_{t}$. Specifically, it is ensured that at least one $l_{t}$ is contained in these subboxes. Using those settings, 6400 subboxes were employed for the training. During the evaluation of testing samples, the entire test sample is reconstructed continuously, without employing cropped reconstruction boxes as is done for training. The velocity components of high-resolution data and low-resolution data used for training and testing were normalized by the global maximum and minimum of the DNS [i.e., $(U_{\textsc{f-dns}},V_{\textsc{f-dns}},W_{\textsc{f-dns}})\in[0,1]$ and $(U_{\textsc{dns}},V_{\textsc{dns}},W_{\textsc{dns}})\in[0,1]$] to improve the network’s performance [32].
@@
The large data set and deep convolutional frameworks (implemented in TensorFlow [33]), required training parallelization; this was done using the Horovod library [34]. Calculations were performed on the Jülich DEEP-EST cluster (DEEP-DAM partition) using four nodes, each with one NVIDIA V100 32GB GPU. This gave an overall speed-up close to a factor of four versus single-GPU training. A mini-batch of eight subboxes per GPU and the ADAM optimizer, with an initial learning rate of $10^{-4}$, were chosen due to memory constraints and previous evidence of the success of this combination [9, 17, 35]. “Mixed-precision” training is common in machine learning for its low computational cost [36] but can be detrimental to predictive accuracy in scientific computing. We use single-precision arithmetic for both weight definition and loss computations, following the results of Hrycej et al. [37]. Details are provided in the Appendix A. Moreover, during the evaluation of the testing samples, CPUs with access to a significantly larger memory pool are employed to overcome memory limitations.
@@
For unsupervised training, LES calculations of forced HIT are conducted using the same DNS solver configuration albeit for coarser grid resolutions (equivalent to the discrete downsampling factor applied to the DNS mesh). The unresolved scales are closed using the dynamic Smagorinsky model [38, 39] [Eq. 11], or their influence is not explicitly modeled (“implicitly modeled” LES). To extrapolate a given model to higher $Re$ (utilizing the Re130 and Re350 data sets for out-of-sample analyses), the filtered input field is rescaled to match the $\Delta/\eta$ ratio used for training. This rescaling procedure guarantees a certain level of generalizability [21].
@@ 3 Neural network architectures and training strategy
Our networks are based on the architecture of Bode et al. [19, 20, 40, 41]. Differing from previous work, our generator network uses additional up-/downsampling layers and dense layers, depicted in Fig. 1; this is inspired by the original ESRGAN architecture [16] and adapted for small-scale turbulence reconstruction [22]. Along with the definition of the two architectures, the training approaches are described next.
@@ 3.1 TSResNet
The TSResNet (Turbulent Super-Resolution Residual Network) is a supervised CNN-type model. It uses only the generator shown in Fig. 1 without the discriminator. The TSResNet is adapted from the original SRResNet [42], which can capture small-scale features when skip connections are included. The generator relies on three-dimensional convolutional layers with Leaky Rectified Linear Units (LReLUs) as activation functions, as they have higher computational efficiency compared to standard rectified linear units (ReLUs) [33]. The Residual-in-Residual Dense Blocks (RRDBs) include residual connections, a set of densely connected-layer blocks (three in the present work), and a residual-in-residual structure [16]. Together, these enable super-resolved data to be generated through a very deep network. This capability is essential to facilitate learning complex transformations and is currently one of the state-of-the-art features of super-resolution networks [43]. Two upsampling layers increase the input spatial resolution by a factor of four in each spatial dimension. Each one doubles the dimensions of the input by nearest-neighbor interpolation (replicating adjacent grid points), followed by a convolution layer to improve the approximation. The upsampling branch (the gray section in Fig. 1) is later modified in Sec. 5.2, where additional upsampling layers are added. The total number of trainable parameters in the TSResNet is approximately $18$ million.
@@ 3.2 PIESRGAN
The training of the GAN architecture is challenging, as the generator and discriminator networks are trained to compete against each other. Finding a convergence point is one of the main challenges of GAN training, and training can suffer oscillations and destabilization in the model’s trainable parameters [13]. To cope with these, the generator is pre-trained in a fully supervised manner (like the TSResNet) by using the original generator loss function proposed in Eq. (6). Subsequently, the pre-trained generator is used to initialize the GAN training. Recent findings demonstrated that the choice of the initial learning rate for both networks is essential for the local convergence of the GAN training, and recommendations from recent literature have been adopted [29, 22].
@@ 3.3 Training strategy
Table 2 lists the architecture, training strategy, and training/testing fields for the specific trained models used in subsequent sections. In the supervised and semisupervised approaches, both LR and DNS field pairs are accessible, and the network utilizes this paired data set in an attempt to learn the mapping function between LR fields and to their corresponding ground-truth counterparts. This methodology is viable only when the data set contains paired data, i.e., both F-DNS and DNS data. In the unsupervised approach, only LR fields (LES data) are available without the corresponding DNS fields, thus learning relies solely on the feedback provided by the discriminator and the physics-informed loss function. The partially supervised approach is a combination of the two strategies, in which F-DNS/DNS pairs and true LES data as well as the discriminator feedback are employed for training.
The TSResNet models were trained for roughly 200 epochs, after which no additional reconstruction improvement was observed, using the Re90 data set and Eq. (6) as the loss function. Statistical convergence required 37 wall-time hours using four NVIDIA Tesla V100 32GB GPUs. The trained TSResNet model parameters were used to initialize the GAN generators. No GAN discriminator pre-training was applied, as this step has recently been demonstrated to have minimal influence [22]. The adversarial training required an additional 11 wall-time hours using the same hardware, i.e., roughly $30\,\%$ more than the TSResNet training time, over an additional 40 epochs. To counteract the influence of the additional training time performed during the GAN training, the TSResNet models were trained for an additional 40 epochs. However, this additional training time did not impact the final reconstructed field and can be considered superfluous.
Model Training strategy Input training – testing fields Filter kernels ($\overline{\Delta}/\Delta_{\textsc{dns}}$) Section
@@
TSResNet (spectral kernel) Supervised S ($4$) – S ($4$) Sec. 4
TSResNet (box kernel) Supervised B ($4$) – B ($4$) Sec. 4
PIESRGAN (spectral kernel) Semisupervised S ($4$) – S ($4$) Sec. 4
PIESRGAN (box kernel) Semisupervised B ($4$) – B ($4$) Sec. 4
@@
PIESRGAN (multiple-kernel-trained) Semisupervised B $\land$ S ($4$) – G ($4$) Sec. 5.1
PIESRGAN (partially unsupervised) Semisupervised and unsupervised B $\land$ S $\land$ LES ($4$) – G ($4$) Sec. 5.1
TSResNet (fixed upsampling) Supervised B ($4$) – B ($8$) Sec. 5.2
PIESRGAN (fixed upsampling) Semisupervised B ($4$) – B ($8$) Sec. 5.2
PIESRGAN ($x\#$) Semisupervised and dynamic upsampling B ($2$land $4$land $8$) – B (##\##) Sec. 5.2
@@
TSResNet Supervised (Re90) B ($4$) – B ($4$) Sec. 6
PIESRGAN Semisupervised (Re90) B ($4$) – B ($4$) Sec. 6
Table 2: List of model configurations, training strategies, and training/testing filter kernels and sizes. The filter kernels “B,” “G,” and “S” abbreviate box, Gaussian, and spectrally sharp kernels, respectively. The $\land$ indicates that the combined set of the input fields is employed in the training, and the ##\## indicates the variable upsampling factor.